OpenStudy (anonymous):

6H2O(l) + 6CO2(g) ----> C6H12O6(s) + 6O2(g) What mass, in grams, of carbon dioxide is needed to react with 3.00 mol of water in this reaction?

OpenStudy (aaronq):

So you have the amount in moles, next you need to build a ratio to relate the moles of water to the moles of carbon dioxide and their respective coefficients. \(\sf\dfrac{moles~of~water}{water's~coefficient}=\dfrac{moles~of~carbon~dioxide}{carbon~dioxide's~coefficient}\)

OpenStudy (aaronq):

Plug your variables in and solve for moles of carbon dioxide. then convert moles to mass (in grams) using: \(\sf moles=\dfrac{mass}{Molar ~mass}\)